OpenStudy (anonymous):

The table below represents the velocity of a car as function of time: time (hour) velocity (miles, hours) x y 0 50 1 52 2 54 3 56 Calculate the average rate of change of the function represented by the table between x = 1 to x = 3 hours, and tell what the average rate represents.

OpenStudy (anonymous):

could someone please help? I have looked through my notes and I don't see much explaining how to do this. I'm just lost with this one right now.

OpenStudy (fortytherapper):

Average Rate of Change is another name for slope, which goes like this: \[\frac{ (y2-y1) }{ (x2-x1) }\]

OpenStudy (anonymous):

yes I understand that! but I don't know what function they are referring to in the first place

OpenStudy (fortytherapper):

They're saying from x=1 to x=3, so in this case it's plug and chug \[\frac{ f(3)-f(1) }{ 3-1 }\] f(3) is the y value at x=3, and the same for f(1)

OpenStudy (anonymous):

so its just 2?

OpenStudy (fortytherapper):

Correct. You can check by looking at the table. Each time X goes up 1, Y goes up 2 y/x = 2/1 = 2

OpenStudy (anonymous):

oh that makes sense! its been a long day lol. Thank you for the help.

OpenStudy (fortytherapper):

Anytime!